OpenStudy (legomyego180):

Simplify:

OpenStudy (legomyego180):

\[\ln[\lim_{z \rightarrow \infty}1+(\frac{ 1 }{ z })^z]+\sin^2(x)+\cos^2(x)=\sum_{n=0}^{\infty}\frac{ \cosh y \sqrt{1-\tanh^2y} }{ 2^n }\]

OpenStudy (sooobored):

1?

OpenStudy (sooobored):

\[\sin^2(x) + \cos^2 (x) = 1\] \[\ln(1) = 0\] \[\sum_{n=0}^{\infty} \frac{1}{2^n} =1\]

OpenStudy (sooobored):

i dont know every trig identity off the top of my head, but 1-tanh^2 is probably a common trig identity which you probably could easily look up

OpenStudy (legomyego180):

Got it. 1+1=2 Thanks.

OpenStudy (anthonylocke):

that is really intimidating math...i'm not that high up yet.

OpenStudy (legomyego180):

Its kind of a joke @AnthonyLocke . Its just a really complicated way of writing 1+1=2